OpenStudy (anonymous):

The formula to find the amount in an account, A, that has an interest rate, r, that compounds n times per year and has a starting balance of P after t years is . If the interest is compounded yearly, then n = 1 and the interest rate, r, represents the annual interest. When the interest is compounded monthly, then n = 12 but r is still the annual interest. a. If you were given the equation , how often is the interest compounded?

OpenStudy (kropot72):

\[\large A=P(1+\frac{r}{n})^{nt}\] where r is expressed as a decimal.

OpenStudy (kropot72):

a) The question states that the interest rate compounds n times per year.

OpenStudy (kropot72):

@Kaelyn78 Are you there?

OpenStudy (anonymous):

here

OpenStudy (anonymous):

thank you

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

b. If you were given the equation , what would the annual interest rate be?

OpenStudy (anonymous):

@kropot72

OpenStudy (kropot72):

As the question states "the interest rate, r, represents the annual interest".

OpenStudy (anonymous):

I don't get it

OpenStudy (kropot72):

"If you were given the equation , what would the annual interest rate be?" The annual interest rate is r in the equation that I posted. r must be expressed as a decimal. If the annual interest rate was 6%, r would be expressed as 0.06.

OpenStudy (kropot72):

@Kaelyn78 Is it any clearer now?

OpenStudy (anonymous):

yes that makes more sense

OpenStudy (anonymous):

c. What would need to change about the equation in part b for it to represent an account that is compounded monthly?

OpenStudy (kropot72):

If an account is compounded monthly, that means it compounds 12 times per year. Which variable in the equation represents the number of times in a year that the account compounds?

OpenStudy (kropot72):

The answer is in the question.

OpenStudy (anonymous):

you need to change r?

OpenStudy (kropot72):

Not really. The variables are A, r, n, P and t. Which one does the question use for "times per year"?

OpenStudy (kropot72):

The question states "When the interest is compounded monthly, then n = 12 but r is still the annual interest."

OpenStudy (anonymous):

so you need to change n?

OpenStudy (kropot72):

Therefore you would replace n by 12 in the equation to represent an account that is compounded monthly.

OpenStudy (kropot72):

Giving: \[\large A=P(1+\frac{r}{12})^{12t}\]

OpenStudy (anonymous):

that makes sense

OpenStudy (anonymous):

d. Use the properties of exponents to rewrite the equation given in part b so that it represents an account that is compounded monthly.

OpenStudy (kropot72):

The required equation was posted above.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

e. What would be the approximate monthly interest rate that is equivalent to the annual interest rate represented in the equation given in part b?

OpenStudy (kropot72):

The approximate monthly interest rate that is equivalent to the annual interest rate represented in the equation given in part b is r/12.

OpenStudy (anonymous):

thank you for all your help